OpenStudy (anonymous):

HELP;WILL GIVE MEDAL TO WHOEVER CAN HELP!!---> What is the solution to the equation 21x + 3 = 14?

OpenStudy (anonymous):

21x+3=14 subtract 3 21x=11 x=11/21

OpenStudy (anonymous):

The equation looks like this: 

OpenStudy (anonymous):

These are the options: x ≈ −2.133 x ≈ −1.846 x ≈ 3.867 x ≈ 4.154

OpenStudy (anonymous):

x=11/21

OpenStudy (anonymous):

That isn't a choice:/

OpenStudy (anonymous):

@ganeshie8 ?